# inkscapepath=svgsubdir

### v1

@@ Abstract
Recent studies have revealed that neural networks learn interpretable algorithms for many simple problems. However, little is known about how these algorithms emerge during training. In this article, we study the training dynamics of a simplified transformer with 2-dimensional embeddings on the problem of modular addition. We observe that embedding vectors tend to organize into two types of structures: grids and circles. We study these structures and explain their emergence as a result of two simple tendencies exhibited by pairs of embeddings: clustering and alignment. We propose explicit formulae for these tendencies as interaction forces between different pairs of embeddings. To show that our formulae can fully account for the emergence of these structures, we construct an equivalent particle simulation where we find that identical structures emerge. We use our insights to discuss the role of weight decay and reveal a new mechanism that links regularization and training dynamics. We also release an interactive demo to support our findings: https://modular-addition.vercel.app/.
@@ 1 Introduction
However, it remains an open problem to provide a similarly high degree of interpretability for the training dynamics that lead to the emergence of these algorithms. Currently, the most successful approach to understanding the learning process is by uncovering hidden progress measures that increase abruptly during training [Barak et al., 2022]. For example, by constructing two progress measures for the problem of modular addition, Nanda et al. [2023] find that training of neural networks on modular addition can be split into three phases: memorization, circuit formation, and cleanup. While insightful, such methods provide only a high-level description of the training process, without offering a clear explanation of the underlying optimization dynamics. Baek et al. [2024] model the embedding vectors as a particle system of “repons” to explain the experimentally observed phase transition between generalization and overfitting, but they do not attemt to provide a complete account of the training dynamics.
The importance of understanding the training dynamics of neural networks is twofold. First, it can contribute to our understanding of neural networks, which in turn allows us to build safer and more interpretable models [Doshi-Velez and Kim, 2017, Olah et al., 2020]. Second, by understanding the training process, we can more easily improve it. We believe that understanding the training dynamics can play a crucial role in the design of new optimizers and architectures, leading to models that are more accurate and compute-efficient.
2 Our contribution
In this article, we study the training dynamics of a simplified transformer with 2-dimensional embeddings on the problem of modular addition. Our contributions are the following.
@@ 2 Our contribution
We show that embedding vectors tend to organize themselves into circular and grid-like structures, as depicted in Figure 1. The emergence of circular and grid-like structures is consistent with the findings of Zhong et al. [2023], Gromov [2023] and Liu et al. [2022]. We show that grids and circles play a key role in the generalization performance of the model. We also show that weight decay plays a crucial role in the emergence of circular structures and in reducing grid imperfections.
@@
We show how grids and circles facilitate accurate classification for the subsequent layers by grouping “pair sums” (outputs of the constant attention). We propose an explanation for the emergence of the grids and circles as a result of two simple phenomena: clustering and alignment. We propose explicit formulae for both phenomena as interaction forces between two different pairs of embeddings.
@@
To prove that our proposed formulae can fully account for the emergence of grids and circles, we construct a particle simulation where particles correspond to embedding vectors and forces correspond to gradients. Forces are computed using only our proposed formulae for clustering and alignment. In this simulation, we show that the particles self-organize into the same structures as the embeddings in the trained transformer: circles, grids, and imperfect grids.
Weight decay
We contribute to the understanding of weight decay by discussing its role in our particular setup. Traditionally, weight decay was understood as a form of $L_{2}$ regularization [Goodfellow et al., 2016]. More recently, it was shown that weight decay plays a much more important role by changing the training dynamics in a desirable way [Zhang et al., 2018]. We show that these two perspectives are in fact complementary: by regularizing one part of the model, we enable the emergence of desirable training dynamics in another part of the model.
We also release an interactive demo to allow the readers to explore the training dynamics and particle simulations for themselves: https://modular-addition.vercel.app/.
@@ 3 Architecture
In this article, we study a simplified single-layer transformer with constant attention. A similar setup has been previously used by Liu et al. [2022], Zhong et al. [2023] and Hassid et al. [2022].
The input to the model consists of two tokens $a$ and $b$ representing the two numbers to be added, where $a,b\in\{0,1,\ldots,N-1\}$. The tokens are embedded into vectors $x_{a}$ and $x_{b}$ using an embedding matrix $E\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. For the scope of this article, we focus on the case when $D=2$ for simplicity and ease of visualization. Then, a constant attention mechanism is applied to the embeddings. This is equivalent to computing the sum of the embedding vectors.
Then we apply a linear layer of size $H$ with weight matrix $W_{h}\in\mathbb{R}^{H\times D}$, bias vector $b_{h}\in\mathbb{R}^{H}$, and ReLU activation. Finally, we apply a linear layer of size $N$ with a weight matrix $W_{o}\in\mathbb{R}^{N\times H}$ and bias vector $b_{o}\in\mathbb{R}^{N}$. The output of the model is the logits of the predicted sum. We don’t use any skip connections or normalization layers.
@@
We train the model in full batch mode using the AdamW optimizer [Loshchilov and Hutter, 2019] with a learning rate of $0.01$ and weight decay between 00 and $1$. We use the cross-entropy loss. Unless specified, we use $N=17$ and $H=32$. The training set consists of 80% of all $N(N+1)/2$ distinct pairs of numbers, chosen randomly. The validation set consists of the remaining 20%.
@@
Weight decay
$\theta_{t+1}\leftarrow\theta_{t}-\lambda\gamma\theta_{t}$ (1)
@@ 4 Grids and circles: the key to successful generalization
We observe that in trained models, the embedding vectors tend to be positioned in arithmetic progressions, forming either grids or circles. We visualize two examples of these structures in Figure 1. The validation accuracy is highest when the embeddings form these structures most clearly. Sometimes, however, the embeddings self-organize into imperfect grids, as shown in Figure 2. In such cases, the validation accuracy is lower, but still higher than when the embeddings are not aligned at all. We present more examples of these structures in Appendix C. To explore the emergence of these structures in more detail, we invite the reader to use our interactive visualizations. In Appendix F, we show that the similar structures emerge in higher-dimensional embedding spaces.
To quantify this effect, we devise two simple algorithms: one to detect circles, and another to quantify the number of grid imperfections in non-circular structures. Circle detection works by checking if all embeddings have about the same distance to the origin. Grid imperfections are defined as the number of pairs of embeddings whose vector sum is not very close to the vector sum of another pair of embeddings with the same modular sum. Both algorithms are explained in detail in the Appendix A.
@@
  • Explanation: We train 100 random initializations for 2000 epochs for various values of weight decay.
    For each value of weight decay, we report the following: the number of circles and their average validation accuracy, the number of non-circles and their average validation accuracy, the average number of grid imperfections for non-circles, and the correlation between the number of grid imperfections and the validation accuracy of non-circles. For circle detection and grid imperfections, see Appendix A.
We run multiple experiments for various values of weight decay and measure the average validation accuracy of circles and non-circles, as well as the correlation between the validation accuracy and the number of grid imperfections of non-circles. We report 95% confidence intervals for the grid imperfections and correlation coefficients [Olivoto et al., 2018]. Results are presented in Table 1. We find that circles consistently have higher validation accuracy than non-circles. For non-circles, we find that the validation accuracy is highly correlated with the number of quantifiable grid imperfections.
The emergence of circles and their high validation accuracy is consistent with the findings of Zhong et al. [2023] and Gromov [2023] who showed that circular structures play an important role in the successful generalization of modular addition. The emergence of grid-like structures is also consistent with the work of Liu et al. [2022] who study similar linear structures in the problem of modular addition.
@@ 5 Co-evolution of embeddings and linear layers
In this section, we will model the training process as the co-evolution of two systems: the embeddings and the subsequent layers. The constant attention mechanism sums the embedding vectors to obtain $x=E_{a}+E_{b}$. The subsequent layers then act as a classifier, trying to classify this value as $c=a+b\ (\mathrm{mod}\ N)$ out of $N$ possible values.
In Figure 3, we visualize the embeddings (left) and the classification function formed by the subsequent layers (right) of a model at the end of training. In the right plot, we also plot the sums of the embedding vectors of all pairs in the training set (for the rest of the article, we will refer to them as simply “pair sums”). Pair sums are colored according to their modular sum. We present more examples in Appendix E. We observe that the grid structure makes the pair sums cluster based on the modular sum, which enables successful classification.
Figure 3: Embedding vectors (left); classifier formed by the subsequent layers
(background, right) and sums of embedding pairs in the training set (markers, right).
In first-order optimization methods, which include the most common optimizers such as SGD and Adam, each parameter is updated according to a gradient that assumes all other parameters are fixed. In practice, this means that each parameter is optimized not only under the constraint of the loss function but also under the constraints of the current state of the other parameters. For our model, this means that the classifier evolves to better classify the pair sums, at the same time that the embeddings evolve to better fit the classifier. We uncover two tendencies that emerge from this dynamic: clustering and alignment.
@@ 5.1 Clustering
As depicted in Figure 3, the pair sums tend to cluster based on their modular sum. Let’s try to understand this phenomenon by considering the interaction between two pair sums $x_{ij}=E_{i}+E_{j}$ and $x_{kl}=E_{k}+E_{l}$, where $i,j,k,l\in\{0,1,\ldots,N-1\}$, and $(i,j)$ and $(k,l)$ are two distinct pairs of numbers in the training set.
Let’s first consider the case when $i+j\not\equiv k+l\ (\mathrm{mod}\ N)$. For simplicity, let’s assume that there are no other pair sums in between $x_{ij}$ and $x_{kl}$. In this case, to classify $x_{ij}$ and $x_{kl}$ correctly, the classifier must place a decision boundary between them. This decision boundary will create a gradient that will push $x_{ij}$ and $x_{kl}$ away from each other. Moreover, the closer $x_{ij}$ and $x_{kl}$ are, the narrower the decision boundary will have to be, resulting in a stronger gradient and a stronger push. In other words, the push will be inversely proportional to the distance between the pair sums. We propose the following formula for the gradient at $x_{ij}$ induced by the pair sum $x_{kl}$:
@@
Now, let’s consider the case when $i+j\equiv k+l\ (\mathrm{mod}\ N)$. Again, for simplicity, let’s assume that there are no other pair sums in between $x_{ij}$ and $x_{kl}$. In this case, $x_{ij}$ and $x_{kl}$ are likely to be inside the same classification region. Both pair sums will be attracted towards the point of maximum classification accuracy, which on average will be the center of the classification region. We will model this case with a constant gradient towards the other pair sum:
@@
To understand why we chose to model the attractive force as a constant (rather that, for example, proportional to the distance between the pair sums), consider the following. The softmax classification boundary resembles a sigmoid: the gradient is maximal at the center of the boundary and decreases with distance. Also, a narrower classification boundary has a stronger gradient at its center but an even weaker gradient farther away. Therefore, the strength of the force applied to a particle will depend on two factors: the width of the classification boundary and the distance to the center of the classification boundary. However, neither of those two factors depends on the distance between the two particles. It might seem that the distance between the particles should be correlated to their distance to the classification boundary, but consider that particles could lie slightly outside of the classification boundary, or even in different classification boundaries altogether.
Note that weight decay plays an important role for both attraction and repulsion. Without weight decay, the weights could get arbitrarily large. We explore the impact of weight decay on the magnitude of the weights in Appendix B. This would lead to classification boundries that are very narrow, producing a clustering force that is very strong in a narrow region, but practically non-existent elsewhere. In other words, without weight decay, the classification function would overfit to the pair sums and the evolution of the embeddings would come to a halt.
@@ 5.2 Alignment
A different picture emerges when we look at a model where the embeddings form a circle. In Figure 4, we visualize the embedding vectors, pair sums, and the classification function of a model with circular embeddings at the end of training. Pair sums are no longer clustered. Rather, they are aligned along lines that pass through the origin. In this arrangement, the classification regions are skewed and start from the origin, resembling a pizza. This model corresponds exactly to the “pizza” algorithm described by Zhong et al. [2023].
Figure 4: Embedding vectors (left); classifier formed by the subsequent layers
(background, right) and sums of embedding pairs in the training set (markers, right).
@@
Weight decay encourages the linear layer to have small weights and biases, which results in linear functions with small slopes and close to the origin. We explore the precise impact of weight decay on the magnitude of the weights and biases in Appendix B. The only aspect that remains completely unconstrained in the direction of the weight vectors. Under these conditions, in order to minimize the training loss, the weight vectors of the linear layer will tend to align with the pair sums. A weight vector aligned with a specific pair sum is maximally useful for the classification layer because it outputs the maximum possible value for that pair sum and the minimum possible value for the other pair sums.
We will model this behavior as follows. For a single pair sum $x_{kl}=E_{k}+E_{l}$, let’s consider a linear function with zero bias ($b=0$) and a weight vector of constant norm that is perfectly aligned with the pair sum ($w=f_{a}\cdot\frac{x_{kl}}{\|x_{kl}\|}$, where $f_{a}$ is a constant). After applying the ReLU activation, the output becomes zero for all points $x$ that oppose the direction of $x_{kl}$ (i.e., $x\cdot x_{kl}<0$). For the other points, the function will have a constant slope of $f_{a}$ in the direction of $x_{kl}$.
Inspired by this simplified model, we propose the following formula for the gradient at $x_{ij}$ produced by the pair sum $x_{kl}$:
@@ 6 Particle simulation
In this section, we show that our proposed model can fully account for the emergence of the grids and circles by constructing a particle simulation and showing that particles converge to the same structures as the embedding vectors in the trained transformer.
@@
We model the training dynamics of the embedding vectors using $N$ particles in a $D$-dimensinal space ($N=17,D=2$). We denote the position of particle $i$ by $x_{i}\in\mathbb{R}^{D}$, $i=1,\ldots,N$. The position $x_{i}$ of particle $i$ corresponds to the embedding vector $E_{i}$ of the number $i$ during training.
We will turn the gradient equations from the previous section into “forces” acting on the $N$ particles. Let $(i,j)$ and $(k,l)$ be two distinct pairs of numbers. We denote the sums of their particle positions as $x_{ij}=x_{i}+x_{j}$ and $x_{kl}=x_{k}+x_{l}$, respectively. The pair $(k,l)$ will induce an equal force on particles $i$ and $j$ (obtained by combining equations 2, 3, and 4):
@@
The force acting on particles $k$ and $l$ is defined analogously. We use $g_{r}=g_{a}=f_{a}=1$. We also weigh the forces so that interactions between pair sums of different modular sum have the same total contribution as those between pair sums of the same modular sum. We initialize the particles randomly according to a normal distribution. At every step of the simulation, we calculate the total force acting on each particle. We then move each particle with a small step in the direction of the total force. We also scale the particle positions to maintain a constant variance and zero mean. We repeat this process for 100 steps. We don’t use any momentum.
We observe that the particles self-organize into the same structures as the embeddings in the trained transformer: circles, grids, and imperfect grids. We visualize a few examples in Figure 5. For more examples, see Appendix D. In Table 2, we present the results of running 100 simulations for various values of $f_{a}$. We find that the relative frequency of circles and the average number of grid imperfections are very similar to the results obtained from the transformer experiments. We also find that increasing $f_{a}$ leads to more circles, which is consistent with our hypothesis that circles emerge as a result of the alignment force.
@@
7 The role of weight decay
Traditionally, weight decay was understood as a form of $L_{2}$ regularization, improving generalization by constraining the network capacity [Goodfellow et al., 2016]. More recently, a new perspective on weight decay has emerged, suggesting that it plays a much more important role during training by changing the training dynamics in a desirable way [Zhang et al., 2018]. In this section, we combine the two perspectives by discussing the impact of weight decay on the training dynamics in our setup.
In Section 4, we observed that weight decay plays an important role in the successful generalization of the trained model. We propose the following explanation: weight decay changes the training dynamics by strengthening the clustering force and by introducing the alignment force. In turn, these forces facilitate the emergence of grids and circles, which are crucial for the generalization performance of the model. Below we discuss exactly how weight decay impacts clustering and alignment.
In the case of clustering, our proposed clustering force (Equations 2 and 3) is highly dependent on a simplified model of the classification function where the decision boundaries are very wide. Weight decay enables this by limiting the magnitude of the weights of the linear layer and output layer, as we show in Appendix B. Without weight decay, classification boundaries could get arbitrarily narrow, leading to a clustering force that is very strong in a narrow region, but practically non-existent elsewhere. In other words, without weight decay, the classifier would overfit to the pair sums and the embeddings would become stuck in a local minimum.
In the case of alignment, weight decay makes the alignment force (Equation 4) possible by limiting the magnitude of the biases of the linear layer, as we show in Appendix B. This forces the ReLU activations of the linear layer to remain very close to the origin, creating the sort of alignment we have seen. Without weight decay, biases could get arbitrarily large and there would not be any specific point around which the alignment could take place.
To the best of our knowledge, this link between regularization and training dynamics has not been previously reported. This opens up interesting new questions for future research, such as providing theoretical proof for this phenomenon or showing how it takes place in other problems and architectures.

### v2

@@ Abstract
Recent studies have revealed that neural networks learn interpretable algorithms for many simple problems. However, little is known about how these algorithms emerge during training. In this article, I study the training dynamics of a small neural network with 2-dimensional embeddings on the problem of modular addition. I observe that embedding vectors tend to organize into two types of structures: grids and circles. I study these structures and explain their emergence as a result of two simple tendencies exhibited by pairs of embeddings: clustering and alignment. I propose explicit formulae for these tendencies as interaction forces between different pairs of embeddings. To show that my formulae can fully account for the emergence of these structures, I construct an equivalent particle simulation where I show that identical structures emerge. I discuss the role of weight decay in my setup and reveal a new mechanism that links regularization and training dynamics. To support my findings, I also release an interactive demo available at https://modular-addition.vercel.app/.
@@ 1 Introduction
However, it remains an open problem to provide a similarly high degree of interpretability for the training dynamics that lead to the emergence of these algorithms. Currently, the most successful approach to understanding the learning process is by uncovering hidden progress measures that increase abruptly during training [Barak et al., 2022]. For example, by constructing two progress measures for the problem of modular addition, Nanda et al. [2023] find that training of neural networks on modular addition can be split into three phases: memorization, circuit formation, and cleanup. While insightful, such methods provide only a high-level description of the training process, without offering a clear explanation of the underlying optimization dynamics. A better understanding of neural networks could lead to AI systems that are more interpretable, more efficient, and more reliable [Doshi-Velez and Kim, 2017, Olah et al., 2020].
2 My contribution
In this article, I study the training dynamics of a small neural network on the problem of modular addition. My architecture consists in a simplified single-layer transformer with constant attention and 2-dimensional embeddings. My contributions are the following.
@@ 2 My contribution
I show that embedding vectors tend to organize themselves into circular and grid-like structures, as depicted in Figure 1. The emergence of circular and grid-like structures is consistent with the findings of Zhong et al. [2023], Gromov [2023] and Liu et al. [2022]. I show that grids and circles play a key role in the generalization performance of the model. I also show that weight decay plays a crucial role in the emergence of circular structures and in reducing grid imperfections.
@@
I show how grids and circles facilitate accurate classification for the subsequent layers by grouping “pair sums” (outputs of the constant attention). I propose an explanation for the emergence of the grids and circles as a result of two simple phenomena: clustering and alignment. I propose explicit formulae for both phenomena as interaction forces between two different pairs of embeddings.
@@
To prove that my proposed formulae can fully account for the emergence of grids and circles, I construct a particle simulation where particles correspond to embedding vectors and forces correspond to gradients. Forces are computed using only my proposed formulae for clustering and alignment. In this simulation, I show that the particles self-organize into the same structures as the embeddings in the trained transformer: circles, grids, and imperfect grids.
I also contribute to the understanding of weight decay by discussing its role in my particular setup in Appendix C. I also release an interactive demo to allow the readers to explore the training dynamics and particle simulations for themselves: https://modular-addition.vercel.app/.
@@ 3 Architecture
In this article, I study a simplified single-layer transformer with constant attention. A similar setup has been previously used by Liu et al. [2022], Zhong et al. [2023] and Hassid et al. [2022].
The input to the model consists of two tokens $a$ and $b$ representing the two numbers to be added, where $a,b\in\{0,1,\ldots,N-1\}$. The tokens are embedded into vectors $x_{a}$ and $x_{b}$ using an embedding matrix $E\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. For the scope of this article, I focus on the case when $D=2$ for simplicity and ease of visualization. Then, a constant attention mechanism is applied to the embeddings. This is equivalent to computing the sum of the embedding vectors.
Then I apply a linear layer of size $H$ with weight matrix $W_{h}\in\mathbb{R}^{H\times D}$, bias vector $b_{h}\in\mathbb{R}^{H}$, and ReLU activation. Finally, I apply a linear layer of size $N$ with a weight matrix $W_{o}\in\mathbb{R}^{N\times H}$ and bias vector $b_{o}\in\mathbb{R}^{N}$. The output of the model is the logits of the predicted sum. I don’t use any skip connections or normalization layers.
@@
I train the model in full batch mode using the Adam optimizer with a learning rate of $0.01$ and decoupled weight decay [Loshchilov and Hutter, 2019] between 00 and $1$. I use the cross-entropy loss. Unless specified, I use $N=17$ and $H=32$. The training set consists of 80% of all $N(N+1)/2$ distinct pairs of numbers, chosen randomly. The validation set consists of the remaining 20%.
@@ 4 Grids and circles: the key to successful generalization
In trained models, the embedding vectors tend to be positioned in arithmetic progressions, forming either grids or circles. I visualize two examples of these structures in Figure 1. The validation accuracy is highest when the embeddings form these structures most clearly. Sometimes, however, the embeddings self-organize into imperfect grids, as shown in Figure 2. In such cases, the validation accuracy is lower, but still higher than when the embeddings are not aligned at all. We present more examples of these structures in Appendix D.
To quantify this effect, I devise two simple algorithms: one to detect circles, and another to quantify the number of grid imperfections in non-circular structures. Both algorithms are explained in detail in the Appendix A.
@@
  • Explanation: I train 100 random initializations for 2000 epochs for various values of weight decay.
    For each value of weight decay, I report the following: the number of circles and their average validation accuracy, the number of non-circles and their average validation accuracy, the average number of grid imperfections for non-circles, and the correlation between the number of grid imperfections and the validation accuracy of non-circles. For circle detection and grid imperfections, see Appendix A.
I run multiple experiments for various values of weight decay and measure the average validation accuracy of circles and non-circles, as well as the correlation between the validation accuracy and the number of grid imperfections of non-circles. I find that circles consistently have higher validation accuracy than non-circles. For non-circles, I find a strong negative correlation between the number of grid imperfections and the validation accuracy.
@@ 5 Co-evolution of embeddings and linear layers
In this section, I will model the training process as the co-evolution of two systems: the embeddings and the subsequent layers. The constant attention mechanism sums the embedding vectors to obtain $x=E_{a}+E_{b}$. The subsequent layers then act as a classifier, trying to classify this value as $c=a+b\ (\mathrm{mod}\ N)$ out of $N$ possible values.
Figure 3: Embedding vectors (left); classifier formed by the combined linear and output
layers (background, right) and sums of embedding pairs in the training set (markers, right).
@@ 5.1 Clustering
As depicted in Figure 3, the pair sums tend to cluster based on their modular sum. I present more examples in Appendix F. Let’s try to understand this phenomenon by considering the interaction between two pair sums $x_{ij}=E_{i}+E_{j}$ and $x_{kl}=E_{k}+E_{l}$, where $i,j,k,l\in\{0,1,\ldots,N-1\}$, and $(i,j)$ and $(k,l)$ are two distinct pairs of numbers in the training set.
Let’s first consider the case when $i+j\not\equiv k+l\ (\mathrm{mod}\ N)$. For simplicity, let’s assume that there are no other pair sums in between $x_{ij}$ and $x_{kl}$. In this case, to classify $x_{ij}$ and $x_{kl}$ correctly, the classifier must place a decision boundary between them. This decision boundary will create a gradient that will push $x_{ij}$ and $x_{kl}$ away from each other. Moreover, the closer $x_{ij}$ and $x_{kl}$ are, the narrower the decision boundary will have to be, resulting in a stronger gradient and a stronger push. In other words, the push will be inversely proportional to the distance between the pair sums. I propose the following formula for the gradient at $x_{ij}$ induced by the pair sum $x_{kl}$:
@@
Now, let’s consider the case when $i+j\equiv k+l\ (\mathrm{mod}\ N)$. Again, for simplicity, let’s assume that there are no other pair sums in between $x_{ij}$ and $x_{kl}$. In this case, $x_{ij}$ and $x_{kl}$ are likely to be inside the same classification region. Both pair sums will be attracted towards the point of maximum classification accuracy, which on average will be the center of the classification region. I will model this case with a constant gradient towards the other pair sum:
@@ 5.2 Alignment
A different picture emerges when we look at a model where the embeddings form a circle. In Figure 4, I visualize the embedding vectors, pair sums, and the classification function of a model with circular embeddings at the end of training. Pair sums are no longer clustered. Rather, they are aligned along lines that pass through the origin. In this arrangement, the classification regions are skewed and start from the origin, resembling a pizza. This model corresponds exactly to the “pizza” algorithm described by Zhong et al. [2023].
Figure 4: Embedding vectors (left); classifier formed by the combined linear and output
layers (background, right) and sums of embedding pairs in the training set (markers, right).
@@
Weight decay encourages the linear layer to have small weights and biases, which results in linear functions with small slopes and close to the origin. We explore the precise impact of weight decay on the magnitude of the weights and biases in Appendix B. The only aspect that remains completely unconstrained in the direction of the weight vectors. A weight vector aligned with a specific pair sum is maximally useful for the classification layer because it outputs the maximum possible value for that pair sum and the minimum possible value for the other pair sums. Thus, under these conditions, in order to minimize the training loss, the weight vectors of the linear layer will tend to align with the pair sums.
We can model this behavior as follows. For a single pair sum $x_{kl}=E_{k}+E_{l}$, let’s consider a linear function with zero bias ($b=0$) and a weight vector of constant norm that is perfectly aligned with the pair sum ($w=f_{a}\cdot\frac{x_{kl}}{\|x_{kl}\|}$, where $f_{a}$ is a constant). After applying the ReLU activation, the output becomes zero for all points $x$ that oppose the direction of $x_{kl}$ (i.e., $x\cdot x_{kl}<0$). For the other points, the function will have a constant slope of $f_{a}$ in the direction of $x_{kl}$.
Inspired by this simplified model, I propose the following formula for the gradient at $x_{ij}$ produced by the pair sum $x_{kl}$:
@@ 6 Particle simulation
In this section, I show that my proposed model can fully account for the emergence of the grids and circles by constructing a particle simulation and showing that particles converge to the same structures as the embedding vectors in the trained transformer.
@@
I model the training dynamics of the embedding vectors using $N$ particles in a $D$-dimensinal space ($N=17,D=2$). I denote the position of particle $i$ by $x_{i}\in\mathbb{R}^{D}$, $i=1,\ldots,N$. The position $x_{i}$ of particle $i$ corresponds to the embedding vector $E_{i}$ of the number $i$ during training.
The gradient equations from the previous section become “forces” acting on the $N$ particles. Let $(i,j)$ and $(k,l)$ be two distinct pairs of numbers. I denote the sums of their particle positions as $x_{ij}=x_{i}+x_{j}$ and $x_{kl}=x_{k}+x_{l}$, respectively. The pair $(k,l)$ will induce an equal force on particles $i$ and $j$ (obtained by combining equations 1, 2, and 3):
@@
The force acting on particles $k$ and $l$ is defined analogously. I use $g_{r}=g_{a}=f_{a}=1$. Forces are weighted so that interactions between pair sums of different modular sum have the same total contribution as those between pair sums of the same modular sum. Particles are initialized randomly according to a normal distribution. At every step of the simulation, the total force acting on each particle is calculated. The, each particle is moved with a small step in the direction of the total force. I also scale the particle positions to maintain a constant variance and zero mean. I repeat this process for 100 steps. No momentum is used.
I observe that the particles self-organize into the same structures as the embeddings in the trained transformer: circles, grids, and imperfect grids. We visualize a few examples in Figure 5. For more examples, see Appendix E. In Table 2, we present the results of running 100 simulations for various values of $f_{a}$. We find that the relative frequency of circles and the average number of grid imperfections are very similar to the results obtained from the transformer experiments. I also find that increasing $f_{a}$ leads to more circles, which is consistent with my hypothesis that circles emerge as a result of the alignment force.